### v2

# MACRO-LLM: LLM-Empowered Multi-Agent Collaborative Reasoning under Spatiotemporal Partial Observability

Abstract
Large Language Model (LLM) agents deployed in complex real-world scenarios increasingly operate as spatially distributed entities. However, this physical dispersion constrains agents to limited local perception and finite temporal horizons. We characterize this bottleneck as spatiotemporal partial observability, where spatial and temporal limitations are fundamentally coupled: resolving spatial conflicts requires temporal reasoning about neighbors’ future actions, while temporal planning requires spatial context beyond local perception. To bridge this gap, we introduce MACRO-LLM, LLM-empowered multi-agent collaborative reasoning under spatiotemporal partial observability. The architecture interleaves spatial and temporal reasoning within each decision cycle via three interdependent modules: (1) the CoProposer mitigates temporal uncertainty by verifying candidate actions via predictive rollouts; (2) the Negotiator overcomes spatial myopia by resolving conflicts through mean-field statistical aggregation, grounded in the CoProposer’s rollout rewards; and (3) the Introspector closes the reasoning loop by analyzing environmental drift and attributing performance changes to refine strategies. Extensive evaluations on two complex long-horizon tasks, cooperative platoon planning and pandemic control, demonstrate that our framework enables robust coordination under spatiotemporal partial observability.
MACRO-LLM: LLM-Empowered Multi-Agent Collaborative Reasoning under Spatiotemporal Partial Observability
Handi Chen, Running Zhao*, Xiuzhe Wu*, Edith C.H. Ngai, The University of Hong Kong, {hdchen, rnzhao}@connect.hku.hk {xzwu, chngai}@eee.hku.hk
footnotetext: *Equal contributionfootnotetext: Corresponding author
1 Introduction
Figure 1: Illustration of spatiotemporal partial observability and the proposed collaborative reasoning framework. (a) Agents are constrained by limited local perception (spatial) and future uncertainty coupled with finite context windows (temporal). (b) Agents mitigate these limitations by exchanging proposals to establish collaborative framework via neighborhood negotiation.
Large Language Model (LLM) agents have demonstrated remarkable human-like capabilities in problem-solving Zhu et al. (2023); Cui et al. (2025). While these agents successfully automate tasks across domains such as healthcare Khasentino et al. (2025) and education Zhao et al. (2025); Shi et al. (2025), single agents often struggle with complexity and scalability in long-horizon scenarios. Inspired by the team collaboration in human societies, researchers are actively exploring systems where multiple LLM agents coordinate to achieve collective goals Li et al. (2023a); Wang et al. (2024). This collective intelligence amplifies individual capabilities, enabling the effective resolution of complex reasoning tasks through coordination and negotiation Zou et al. (2025). However, most existing frameworks assume shared global context or unrestricted communication among agents. As LLM agents increasingly operate in physical-world scenarios, coordination under limited local perception remains underexplored.
In such settings, agents operate as distributed entities Chen et al. (2025), which inherently impose spatiotemporal partial observability. Specifically, physical dispersion constrains agents to limited local perception, while the LLM’s fixed context window restricts their historical temporal information. This challenge is widespread: in pandemic control, districts must coordinate containment policies with only local infection data; in traffic coordination, vehicles optimize flow under communication limits and privacy constraints. Therefore, addressing the partial observability stemming from real-world distribution is pivotal for enhancing collaborative efficiency and effectiveness.
Specifically, partial observability manifests along two dimensions: spatial and temporal, as illustrated in Fig. 1(a). Spatially, constrained by communication limits and environmental complexity, each agent perceives only a fraction of the global information, preventing it from inferring the full context of the collective task. Temporally, the stochastic environment and unpredictable behaviors of others make long-term consequences difficult to predict, while finite historical context further hinders alignment of short-term decisions with long-term objectives. Together, these constraints result in suboptimal or even conflicting actions that degrade collective efficiency and effectiveness.
Existing LLM-based Multi-Agent Systems (MAS) aim to expand their spatial perception through inter-agent communication, as shown in Fig. 1 (b). This is typically implemented via hierarchical structures that aggregate information to a central node Yu et al. (2024); Estornell et al. (2025) or fully-connected graphs that construct a pseudo-global view Chan et al. (2023). However, these methods often incur communication bottlenecks at central nodes, limiting their scalability in large-scale, dynamically evolving environments. Temporally, while memory modules help mitigate limited historical observations Park et al. (2023); Zhang et al. (2024), the inherent unpredictability of future states remains unaddressed. This uncertainty makes it difficult for agents to align immediate actions with long-term goals and capture cascading temporal effects. Prior to LLM-based MAS, Multi-Agent Reinforcement Learning (MARL) provides a traditional framework for partial observability but suffers from poor generalization and excessive training costs Huh and Mohapatra (2023), limiting its feasibility for complex real-world applications.
In this paper, we propose MACRO-LLM, an LLM-empowered multi-agent collaborative reasoning under spatiotemporal partial observability, generalizable across diverse domains. A key insight is that spatial and temporal partial observability are coupled: resolving spatial conflicts requires predicting how neighbors will act in the future, while planning over time requires spatial context beyond local perception. Existing approaches address these dimensions in isolation. MACRO-LLM bridges this gap through a negotiation-centric architecture where spatial and temporal reasoning are interleaved within each decision cycle via three interdependent modules: the CoProposer, which generates collaborative proposals validated via predictive rollouts to mitigate temporal uncertainty; the Negotiator, which resolves conflicts through mean-field statistical aggregation of unobservable agents to overcome spatial myopia; and the Introspector, which refines strategies via scenario-aware adaptive reflection to ensure continuous adaptation. Our main contributions are summarized as follows:
Insightful.
We explicitly decompose partial observability into spatial and temporal dimensions and address them jointly through a unified decentralized negotiation framework that couples semantic reasoning with mean-field statistical features, enabling robust long-horizon coordination without central coordinators or global views.
Technical.
We develop MACRO-LLM with three interdependent modules forming a closed reasoning loop: the Negotiator’s confidence assessment consumes the CoProposer’s rollout rewards, and the Introspector’s revision signals reshape the strategies for subsequent cycles. This interdependence constitutes the core contribution beyond the individual techniques each module builds upon.
Experimental.
We evaluate on two complementary domains, Cooperative Platoon Planning (CPP) and Pandemic Coordination (PC), covering continuous and discrete actions, linear and graph topologies, fine-grained and long-horizon planning. Results demonstrate MACRO-LLM’s superior coordination performance, scalability, and robustness across six LLM backbones.
2 Related Works
LLM-based Multi-Agent Systems.
Existing LLM-based MAS have evolved from general collaboration frameworks Li et al. (2023a); Wu et al. (2023) to structured systems with persistent memory and reasoning capabilities Hong et al. (2023); Wang et al. (2023). To address spatial partial observability, recent works adopt either hierarchical aggregation, which centralizes information flow Yu et al. (2024); Estornell et al. (2025), or belief-based inference, which deduces hidden states via reasoning chains Li et al. (2023b); Davidson et al. (2024). However, these approaches often rely on ideal communication bandwidth or static topologies, failing to account for the dynamic coupling of spatial dispersion and temporal uncertainty. Additionally, swarm-based optimization methods Feng et al. (2024); Zhang et al. (2025) evolve agent configurations through offline iterative search, yielding static configurations that cannot adapt to runtime dynamics and lack coordination mechanisms for resolving conflicts under partial observability (see Appendix A). MACRO-LLM fills this gap by enabling decentralized spatiotemporal reasoning without reliance on central nodes or global views.
Multi-Agent Reinforcement Learning.
Traditional approaches address partial observability via the Dec-POMDP framework Oliehoek et al. (2016); Koops et al. (2024), often utilizing Centralized Training with Decentralized Execution (CTDE) Kia et al. (2024); Bernstein et al. (2005) or Mean-Field approximations Yang et al. (2018); Ganapathi Subramanian et al. (2020) to enhance scalability. While algorithms like MAPPO Schulman et al. (2017), DMPO Ma et al. (2024), and communication-aware methods like IC3Net Singh et al. (2018) improve stability, they remain computationally intensive and environment-specific. They require costly retraining when transferring to new scenarios due to poor generalization Huh and Mohapatra (2023); Oroojlooy and Hajinezhad (2023). In contrast, our framework leverages LLMs to achieve zero-shot generalization without parameter updates. Due to space constraints, we provide an extended review in Appendix A.
3 MACRO-LLM Framework
Figure 2: Architecture of MACRO-LLM for agent $n$. The framework comprises three synergistic modules: (1) the CoProposer generates proposals; (2) the Negotiator handles conflict resolution and spatial strategy updates; and (3) the Introspector performs strategy refinement. The CoProposer and Negotiator form a negotiation loop to process observations at time $t$ and output coordinated actions, leading to the next observation at $t+1$.
MACRO-LLM establishes a decentralized negotiation framework for MAS operating under spatiotemporal partial observability. As illustrated in Fig. 2, the architecture of agent $n$ comprises three synergistic modules: (1) CoProposer, for collaborative proposal generation; (2) Negotiator, for conflict resolution and spatial strategy updates; and (3) Introspector, for temporal strategy refinement via scenario-aware adaptive reflection.
3.1 System Overview
We model the MAS as a peer-to-peer graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes agents and $\mathcal{E}$ communication links. For agent $n\in\mathcal{V}$, its neighborhood is $\mathcal{N}_{n}=\{m\in\mathcal{V}|(n,m)\in\mathcal{E}\}$. The system operates without a central coordinator. At initialization, each agent receives a system prompt defining the task, action space, state dynamics, and reward criteria. Architecturally, spatial and temporal reasoning are not separable. Evaluating an action requires predicting neighbors’ responses (temporal), but those predictions depend on unobservable states estimable only via mean-field spatial aggregation. MACRO-LLM operationalizes this dependency by interleaving the CoProposer and Negotiator within each negotiation round, with the Introspector providing cross-cycle feedback.
3.2 CoProposer
The CoProposer generates collaborative proposals $P_{n}^{t}$ comprising actions for agent $n$ and its observable neighbors $\mathcal{N}_{n}$ by integrating long-term objectives with immediate spatial constraints.
Temporal Strategy Formulation.
To align short-term decisions with global objectives despite incomplete information, we introduce a high-level temporal strategy. Agent $n$ derives this strategy based on its current observation $o_{n}^{t}$, action space $\mathcal{A}_{n}$, and global objectives $\mathcal{J}^{global}$. This produces a macroscopic plan $\Pi_{n,t}^{\text{temp}}=\text{LLM}(o_{n}^{t},\mathcal{A}_{n},\mathcal{J}^{global})$, guiding step-by-step progression toward the goal.
Spatial Strategy Formulation.
To incorporate spatial context, agent $n$ infers potential cascading effects originating from unobservable agents. By analyzing its local observation $o_{n}^{t}$ and the network topology $\mathcal{G}$, the agent forecasts multi-agent dynamics to formulate a spatial strategy $\Pi_{n,t}^{\text{spatial}}$.
Proposal Generation via Rollout-Simulated Verification.
Leveraging current spatial and temporal strategies, agent $n$ first generates an initial proposal $P_{n}^{t,0}=(o_{n}^{t},a_{n}^{t,0},\{a_{n\rightarrow m}^{t,0}\})$ by identifying the most advantageous action for itself and inferring cooperative actions for its neighbors $\mathcal{N}_{n}$. Due to future uncertainty, this proposal requires validation through rollout-simulated verification, performed via LLM inference without simulator access. Agent $n$ predicts $\hat{o}_{n}^{t+1,0}$ by reasoning over domain dynamics, $\Pi_{n,t}^{\text{temp}}$, and $\Pi_{n,t}^{\text{spatial}}$, then assesses whether the predicted state constitutes an improvement. This process iterates over $k$ steps, constituting a rollout.
Rollout: In our framework, a rollout is a multi-step LLM-based projection from time $t$ to $t+k$ that infers agent $n$’s future states and actions under its current strategy. The sequence of predicted observations, actions, and rewards across these $k$ steps forms a trajectory, denoted as $\tau_{n}=\Big((o_{n}^{t},a_{n}^{t},R_{n}^{t}),\dots,(o_{n}^{t+k},a_{n}^{t+k},R_{n}^{t+k})\Big)$.
To evaluate a rollout, we employ a progressive constraint relaxation mechanism. For the immediate step ($t$), the action must strictly satisfy all environmental constraints to ensure safety. For subsequent steps ($t+1$ to $t+k$), considering increasing uncertainty, criteria are relaxed to minimum viability constraints. This ensures the action $a_{n}^{t}$ is safe while conducive to future positive outcomes. The estimated cumulative reward is formulated as:
$\mathcal{R}(\tau_{n})=\sum_{\iota=0}^{k}\gamma^{\iota}R_{n}(o_{n}^{t+\iota},a_{n}^{t+\iota}).$ (1)
To reduce computational complexity over long horizons, reward signals beyond $t+1$ are simplified to a binary indicator ($r=1$ if constraints are met, else 0). If a proposal violates constraints during the rollout, agent $n$ revises its proposed action until a valid proposal is found or the maximum attempt limit is reached. The candidate with the highest $\mathcal{R}(\tau_{n})$ is selected and broadcast to neighbors. The pseudocode is provided in Appendix B.
3.3 Negotiator
In decentralized environments, agents generate proposals concurrently, inevitably introducing conflicts that necessitate a negotiation mechanism.
Neighborhood State Aggregation via Mean-Field Approximation.
To extend agents’ perception beyond observable limits, we complement semantic reasoning with statistical state features. In complex topologies, exchanging raw observations is inefficient due to bandwidth constraints. To address this, we introduce a protocol based on mean-field approximation to transmit aggregated statistical features.
The core idea of mean-field theory approximates complex pairwise interactions between an individual and all other agents in the environment by the interaction between the individual and an aggregated “mean field” effect Lasry and Lions (2007). Accordingly, we represent the neighborhood influence through weighted statistics. For an agent $n$, the state $s_{m}^{t}$ of a neighbor $m$ is decomposed into a local mean field and a fluctuation term:
$s_{m}^{t}=\mu_{n}+\delta s_{n\rightarrow m}^{t},\quad\mu_{n}=\frac{\sum_{m\in\mathcal{N}_{n}}w_{n,m}s_{m}^{t}}{\sum_{m\in\mathcal{N}_{n}}w_{n,m}},$ (2)
where $\mu_{n}$ represents the weighted mean of the neighborhood’s spatial features. The weighted variance is defined as:
$\sigma_{n}^{2}=\frac{\sum_{m\in\mathcal{N}_{n}}w_{n,m}(s_{m}^{t}-\mu_{n})^{2}}{\sum_{m\in\mathcal{N}_{n}}w_{n,m}}.$ (3)
Here, weights $w_{n,m}>0$ are determined by domain-specific relational factors, such as the spatial distance, relative speed between agents. During negotiation, agents transmit these statistical features instead of raw data to characterize unobservable contexts. Once agent $n$ receives statistical features $(\mu_{m},\sigma_{m}^{2},W_{m})$ from a neighbor $m$, it updates the aggregated features incorporating its own state $s_{n}^{t}$ (with weight $w_{n}$) using the weighted Welford algorithm Welford (1962). Since each neighbor already aggregates its own neighborhood’s states into the transmitted statistics, this enables agent $n$ to incorporate spatial insights without requiring access to the raw states of all agents. The update equation is provided in Appendix F.
Eq. 1 defines the estimated cumulative reward $\mathcal{R}(\tau_{n})$ for agent $n$’s rollout trajectory $\tau_{n}$. The trajectory $\tau_{n}$ depends on the neighborhood states $\mathbf{s}_{\mathcal{N}_{n}}^{t}=(s_{1}^{t},\ldots,s_{|\mathcal{N}_{n}|}^{t})$, which influence the agent’s spatial strategy $\Pi_{n,t}^{\text{spatial}}$ (Sec. 3.2). Eq. 1 can be rewritten as $\mathcal{R}_{n}(\mathbf{s}_{\mathcal{N}_{n}}^{t})\;\triangleq\;\mathcal{R}(\tau_{n})$ to expose the neighborhood dependence. Therefore, the error is denoted as $\left|\mathcal{R}_{n}(\mathbf{s}_{\mathcal{N}_{n}}^{t})-\mathcal{R}_{n}(\boldsymbol{\mu}_{n})\right|$ i.e., the gap between the rollout reward computed with full neighborhood information $\mathbf{s}_{\mathcal{N}_{n}}^{t}$ and with the mean-field approximation $\boldsymbol{\mu}_{n}=(\mu_{n},\ldots,\mu_{n})$.
To justify transmitting $\sigma^{2}_{n}$ as an approximation quality indicator, we invoke Eqs. (7)(8) and Appendix B of Yang et al. (2018). In their formulation, pairwise Q-function interactions are approximated by the interaction with a mean agent, with the approximation error controlled by the variance of deviations from the mean. We apply the principle to analyze the theoretical properties of our mean-field aggregation, yielding the error bound:
$\left|\mathcal{R}_{n}(s^{t}_{\mathcal{N}_{n}})-\mathcal{R}_{n}(\mu_{n})\right|\leq\frac{L_{2}}{2w_{\min}}\cdot\sigma^{2}_{n},$ (4)
where $L_{2}$ is a smoothness constant determined by the continuity of the underlying domain dynamics and the boundedness of the state-action spaces, which hold broadly in physically grounded environments with continuous dynamics and finite horizons (a detailed discussion is provided in Appendix C). $\sigma^{2}_{n}$ controls the approximation error bound. In many topologies, agent $n$ acts as a cut vertex connecting disjoint subgroups (e.g., predecessor and follower in the CPP linear topology). Since these subgroups lack direct communication, their state distributions naturally diverge, making a single $\mu_{n}$ a poor summary. We prove that computing separate statistics per subgroup yields a provably tighter bound:
Theorem 1 (Tighter Bound via Topology-Aware Partitioning).
With the general error bound detailed in Eq. 4, computing separate weighted statistics per subgroup yields a tighter bound $\left|\mathcal{R}_{n}(\mathbf{s}_{\mathcal{N}_{n}}^{t})-\mathcal{R}_{n}(\boldsymbol{\mu}_{n}^{(1:P)})\right|\leq\frac{L_{2}}{2\,w_{\min}}\sum_{p=1}^{P}|\mathcal{N}_{n}^{(p)}|\cdot\sigma_{n}^{2,(p)}\leq\frac{L_{2}}{2\,w_{\min}}\cdot|\mathcal{N}_{n}|\cdot\sigma_{n}^{2}$.
The proof is provided in Appendix D. Both bounds are controlled by $\sigma_{n}^{2}$. Thus $\sigma_{n}^{2}$ serves a dual role: (1) error indicator: small $\sigma_{n}^{2}$ guarantees accurate mean-field approximation; (2) uncertainty signal: large $\sigma_{n}^{2}$ prompts conservative strategies. During negotiation, agents progressively align their strategies, reducing $\sigma_{n}^{2}$ and tightening both bounds over successive rounds.
Proposal Confidence Assessment and Regeneration.
To evaluate proposal reliability, agent $n$ computes the pairwise differences between its own proposal $P_{n}^{t}$ and received neighbor proposals $\{P_{m}^{t}\}_{m\in\mathcal{N}_{n}}$. Local consensus between agents $n$ and $m$ is defined as $|P_{n}^{(q)}-P_{m}^{(q)}|<\delta$, where $\delta=\epsilon\cdot(a_{\max}-a_{\min})$ is proportional to the action space range for continuous actions, and $\delta=0$ for discrete action spaces requiring exact agreement. If consensus is not met, the agent analyzes discrepancies, updates its spatial strategy by incorporating the received mean-field states, and integrates weighted neighbor proposals to regenerate a refined candidate, which then undergoes rollout verification as detailed in Sec. 3.2.
Multi-Round Negotiation.
A complete iteration of the CoProposer (Sec. 3.2) and Negotiator (Sec. 3.3) constitutes a negotiation round. The primary purpose of multi-round negotiation is information propagation. Each round extends agents’ effective perception by one additional communication hop, with $r$ rounds providing coverage of $r+1$ hops. Local consensus serves as an early stopping criterion. Upon completion, agent $n$ produces a final decision by weighting each neighbor’s proposal by the confidence scores from proposal evaluation, selecting the highest-confidence action (or weighted combination for continuous actions). We provide Theorem 2, which establishes that proposal quality is monotonically non-decreasing across rounds.
Theorem 2 (Properties of the Negotiation Loop).
Let $\mathcal{G}$ be a connected graph and $r$ be the maximum negotiation rounds. The negotiation loop satisfies: (1) Bounded Termination: The negotiation terminates in at most $r$ rounds, either via early stopping when local consensus is reached ($\|P_{n}-P_{m}\|<\delta,\ \forall m\in\mathcal{N}_{n}$) or via the maximum round limit. (2) Monotonic Non-Decrease of Proposal Quality: $\mathcal{R}(\tau_{n}^{(\ell+1)})\geq\mathcal{R}(\tau_{n}^{(\ell)}),\quad\forall\,\ell=0,1,\ldots,r-1$.
The proof can be found in Appendix E. Regardless of whether agents reach consensus, the negotiation process terminates within $r$ rounds, and proposal quality is guaranteed to be monotonically non-decreasing across rounds, exhibiting a convergent trend toward improved coordination.
3.4 Introspector
Due to spatiotemporal partial observability, an agent’s state inference is often compromised by dynamic environmental shifts and the unpredictable behaviors of others. Consequently, if an agent observes a reward decline between time steps, it initiates a self-reflection mechanism to update strategy. We model this process as scenario-aware adaptive reflection, which determines how much to revise based on environmental drift and what to revise through attribution-based reasoning.
Revision Intensity Estimation.
We first quantify the magnitude of environmental shifts to determine the revision intensity. Let $o_{n}^{t}$ denote the vectorized observation of agent $n$ at time $t$. We construct a transition vector $\Gamma_{n}^{t}$ by concatenating the min-max normalized observation difference and the action vector, defined as $\Gamma_{n}^{t}=(\Delta{\tilde{o}_{n}^{t}},\tilde{a}_{n}^{t})$, where $\Delta{o_{n}^{t}}=o_{n}^{t+1}-o_{n}^{t}$.
To evaluate scene consistency, we compute the cosine similarity between two consecutive transitions, $\Gamma_{n}^{t-1}$ and $\Gamma_{n}^{t}$. The drift intensity ($\lambda_{\text{drift}}$) is derived from scene discontinuity:
$\lambda_{\text{drift}}=1-\frac{\Gamma_{n}^{t-1}\cdot\Gamma_{n}^{t}}{\|\Gamma_{n}^{t-1}\|\cdot\|\Gamma_{n}^{t}\|}.$ (5)
Higher $\lambda_{\text{drift}}$ indicates greater environmental drift requiring aggressive revision, while lower values permit only fine-grained refinements. Combined with the reward-decline-based trigger condition to ensure genuine performance degradation.
Revision Direction Identification.
We generate a semantic revision signal as a natural language description of the “direction” for strategic improvement. The LLM analyzes observation logs and negotiation records $\mathbf{N}_{t}$ to identify which transitions deviated from the temporal strategy and to assess the impacts from neighbors. The revision signal $g_{n,t}^{\text{semantic}}$ is generated as textual rules that characterize the associations between actions, state changes, and negative reward outcomes, providing diagnostic guidance for strategy refinement.
Strategy Update.
Finally, we update the temporal strategy using the revision signal and drift intensity. A concrete prompt template for strategy update with revision intensity and direction is provided in Appendix Q.3. The update is formulated as $\Pi_{n,t+1}^{\text{temp}}\leftarrow\text{LLM}\left(\Pi_{n,t}^{\text{temp}},\ g_{n,t}^{\text{semantic}},\ \lambda_{\text{drift}}\right)$. The drift intensity $\lambda_{\text{drift}}$ modulates the revision scope: a high $\lambda_{\text{drift}}$ indicates substantial environmental drift, instructing the LLM to aggressively revise the strategy according to $g_{n,t}^{\text{semantic}}$; a low $\lambda_{\text{drift}}$ confines revision to fine-grained adjustments, preserving the core of the existing strategy $\Pi_{n,t}^{\text{temp}}$.
4 Experiments
4.1 Experimental Settings
Table 1: Performance comparison on the CPP task across Catch-up and Slow-down scenarios. Metrics include RMSE and SD for Headway (H) and Velocity (V). Bold and underline indicate best and second-best results, respectively.
[…]
Methods RMSE-H ($\downarrow$) RMSE-V ($\downarrow$) SD-H ($\downarrow$) SD-V ($\downarrow$)
ToM-Belief 5.514 4.409 4.002 3.130
ChatEval 5.495 2.113 2.947 1.241
LAMEN 1.891 0.648 3.320 0.601
DPPO 1.603 0.978 0.994 0.474
DMPO 4.765 0.534 3.025 1.349
IC3Net 8.687 4.205 7.232 3.399
[…]
Methods RMSE-H ($\downarrow$) RMSE-V ($\downarrow$) SD-H ($\downarrow$) SD-V ($\downarrow$)
ToM-Belief 0.641 4.681 0.225 0.480
ChatEval 3.062 5.708 0.778 1.016
LAMEN 0.923 3.385 0.635 0.750
DPPO 1.720 3.121 1.410 1.176
DMPO 2.226 3.059 1.742 1.189
IC3Net 7.877 4.678 5.207 2.319
[…]
MACRO-X: DS denotes DeepSeek-V3.1, Llama denotes
Llama-3.3-70B-Instruct, and GPT denotes GPT-4o.
Task Configurations.
We evaluate MACRO-LLM across two complementary tasks where coordination outcomes are measured through domain-specific physical metrics, ensuring that performance gains reflect genuine coordination ability rather than language fluency. (1) Cooperative Platoon Planning (CPP) Chu et al. (2020): a multi-agent vehicle platooning task (default N=8) with dynamic “Catch-up” and “Slow-down” scenarios. (2) Pandemic Control (PC) Kompella* et al. (2020): a strategic planning task where agents select regulation level to mitigate virus spread across three urban topologies (Helsinki, Hong Kong, and New York). CPP and PC tasks span continuous and discrete action spaces, linear and mixed topologies, and fine-grained and long-horizon decision-making, respectively. Refer to Appendix G for simulator configurations and city topologies.
Implementation Details.
We implement MACRO-LLM as a backbone-agnostic framework and evaluate it across six LLM backbones: GPT-4o, GPT-4o-mini, DeepSeek-V3.1, Llama-3.3-70B-Instruct, Mistral-Small-3.2-24B-Instruct, and Qwen3-flash. Tables 1 and 2 report three representative backbones; the full generalizability analysis is in Appendix K. Scalability, ablation, and robustness analyses use GPT-4o as the default. Framework and task configurations are detailed in Appendix G. Results for MACRO-LLM are averaged over 5 independent runs; variance analysis confirming low standard deviations across all metrics is provided in Appendix N. A case study on hierarchical deployment with different LLM invocation frequency is in Appendix I.
Table 2: Performance comparison on the PC task across Helsinki, Hong Kong, and New York topologies. Metrics include Normalized Infection (I_n), Peak Infection (PI_n), Normalized Death (D_n), and Pandemic Duration (PD). Bold and underline indicate best and second-best results, respectively.
Methods Helsinki HongKong NewYork
I_n ($\downarrow$) PI_n ($\downarrow$) D_n ($\downarrow$) PD ($\downarrow$) I_n ($\downarrow$) PI_n ($\downarrow$) D_n ($\downarrow$) PD ($\downarrow$) I_n ($\downarrow$) PI_n ($\downarrow$) D_n ($\downarrow$) PD ($\downarrow$)
ToM-Belief 0.184 0.042 0.002 92 0.239 0.028 0.003 120 0.130 0.023 0.001 60
ChatEval 0.010 0.008 0.000 42 0.009 0.003 0.001 24 0.290 0.043 0.002 83
LAMEN 0.010 0.008 0.002 17 0.012 0.005 0.000 23 0.164 0.025 0.002 52
DPPO 0.426 0.110 0.007 62 0.370 0.083 0.004 83 0.923 0.363 0.009 95
DMPO 0.010 0.008 0.001 13 0.271 0.057 0.003 59 0.332 0.109 0.002 53
IC3Net 0.651 0.186 0.009 69 0.259 0.067 0.003 52 0.789 0.284 0.007 93
[…]
MACRO-X: DS denotes DeepSeek-V3.1, Llama denotes Llama-3.3-70B-Instruct, and GPT denotes GPT-4o.
Baselines.
We benchmark MACRO-LLM against representative methods from two paradigms: (1) LLM-based MAS baselines, which share the same backbone (GPT-4o) and environmental interfaces to provide controlled comparisons: ToM-Belief Li et al. (2023b), ChatEval Chan et al. (2023), and LAMEN Davidson et al. (2024); and (2) MARL baselines, which serve as task-specialized performance references trained on each topology: DPPO, DMPO Ma et al. (2024), and IC3Net Singh et al. (2018). Details on baseline implementations, prompt adaptations, and the division between LLM reasoning and external numerical modules are provided in Appendix H.3.
Evaluation Metrics.
For the CPP task, we evaluate platoon stability via the Root Mean Square Error (RMSE) and Standard Deviation (SD) of both velocity and headway distance (denoted as RMSE-V/H and SD-V/H). For the PC task, we assess containment efficacy through Normalized Infection (I_n), Peak Infection (PI_n), Normalized Death (D_n) and total Pandemic Duration (PD). Metric definitions are detailed in Appendix J.
4.2 Performance of Cooperative Platoon Planning Task
As shown in Table 1, MACRO-LLM consistently outperforms both MARL and LLM-based baselines across multiple backbones. In the Catch-up scenario, all three MACRO-LLM variants achieve the top RMSE-H and SD-H results. The SD-H of MACRO-GPT ($0.409\pm 0.034$, Appendix N) is 87.7% lower than that of LAMEN and 58.9% lower than that of DPPO, while even MACRO-DS attains lower SD-H than all baselines. In the Slow-down scenario, MACRO-LLM prioritizes headway safety over aggressive velocity matching. MACRO-GPT achieves the lowest SD-H ($0.128\pm 0.054$, Appendix N), a 43.1% improvement over ToM-Belief, while maintaining competitive velocity errors. The consistent advantage across three diverse backbones confirms that the performance gains stem from the framework’s negotiation mechanism rather than any single model’s capability.
Figure 3: Scalability analysis of MACRO-LLM on the CPP task for number of agents $N\in\{8,16,24,32\}$.
4.3 Performance on Pandemic Control Task.
Table 2 presents the comparative results across three city topologies. MACRO-LLM consistently achieves the lowest infection rates and shortest pandemic durations across all settings. In New York topology, all three MACRO-LLM variants reduce cumulative infections by over 99% compared to DPPO while shortening the pandemic duration from 95 to as few as 11 days (MACRO-DS). In Hong Kong, MACRO-Llama reduces PD to 16 days, outperforming LAMEN (23 days) at comparable infection levels. MACRO-LLM achieves both low infections and short durations simultaneously. This consistent advantage across three diverse backbones confirms that the coordination gains stem from the framework’s negotiation mechanism rather than any single model’s capability.
4.4 Analysis of Scalability
To investigate scalability, we extend the CPP Catch-up scenario to fleet sizes $N\in\{8,16,24,32\}$. As illustrated in Fig. 3, MACRO-LLM maintains consistent robustness as the network quadruples. Oscillation amplitudes remain bounded and all agents converge to the target equilibrium within 40s regardless of fleet size. Aggregated metrics (Appendix L, Table 6) reveal two complementary trends: RMSE-H decreases from 1.237 (NN=8) to 0.728 (NN=32), as the growing proportion of already-stabilized agents dilutes the transient deviations of those still responding to propagation. Meanwhile, SD-H increases moderately from 0.409 to 0.433, reflecting the inherently longer synchronization lag in larger platoons. Importantly, this increase is bounded. At NN=32, SD-H remains 56.4% lower than that of DPPO (NN=8), confirming that MACRO-LLM compresses large-scale interactions while ensuring coordination quality. Further scalability analyses are provided in Appendix L.
4.5 Ablation Study
Figure 4: Ablation of MACRO-LLM—impact of each module on CPP performance.
Figure 5: Ablation of MACRO-LLM—impact of each module on PC performance.
We validate each module’s contribution in Figs. 4 and 5. Ablating the Negotiator degrades stability metrics drastically. Without the Introspector, agents fail to suppress oscillations over time. Notably, the ablations reveal cross-module dependencies: removing the CoProposer disables not only temporal verification but also the Negotiator’s confidence assessment without rollout rewards. It explains the largest SD-H degradation among all variants. Quantitative analyses are in Appendix M.
5 Conclusion
In this work, we present MACRO-LLM, a novel framework enabling agents to reason and coordinate through localized negotiations under spatiotemporal partial observability. We explicitly address the coordination challenge along spatial and temporal dimensions, addressing them through three interdependent modules: (1) a CoProposer to generate collaborative actions and mitigate future uncertainty via predictive rollouts; (2) a Negotiator that resolves spatial conflicts using mean-field statistical estimations; and (3) an Introspector that facilitates strategic refinement via scenario-aware adaptive reflection. Extensive experiments on CPP and PC tasks demonstrate MACRO-LLM’s capability to align local actions with long-term global objectives, achieving superior collaborative efficiency compared to representative baselines. Ultimately, our work highlights the critical role of explicitly modeling spatiotemporal dynamics in scaling LLM-based MAS.
Limitations
While MACRO-LLM demonstrates robust coordination capabilities, we acknowledge two primary limitations. First, as a strategic-level reasoning framework, MACRO-LLM’s iterative negotiation mechanism prioritizes decision quality over inference speed. The per-step latency is dominated by API network overhead of remote LLM calls, which is orthogonal to the framework’s reasoning complexity and can be mitigated through two composable strategies already validated in our experiments: (1) hierarchical deployment, where MACRO-LLM issues strategic directives every $\Delta$ steps while a lightweight controller maintains millisecond-level safety guarantees. Empirical results confirm that even with 90% fewer LLM invocations, coordination quality surpasses all baselines (Appendix I, Table 4); and (2) local deployment of lightweight models such as Qwen-flash and Mistral-small-3.2-24B-Instruct, which eliminates API latency entirely while maintaining competitive performance (Appendix K). Future work will explore tighter co-design between the strategic layer and domain-specific execution layers to further optimize the reasoning-latency balance. Second, relying on foundation models requires additional context injection for domain adaptation, which increases token consumption. Future work will explore online self-evolution mechanisms, such as experience-driven prompt distillation and adaptive strategy caching, that enable agents to progressively internalize domain-specific knowledge through interaction, reducing context dependency while improving cross-domain adaptability.
Appendix Table of Contents
Appendix A Extended Related Works
Due to space constraints, we provide a comprehensive review of the literature in this appendix.
Agent Planning with LLMs.
There has been a recent surge in research on LLM-based agent reasoning and planning Huang et al. (2024); Yuksekgonul et al. (2025). Techniques such as Chain-of-Thought (CoT) Wei et al. (2022), ReAct Yao et al. (2023), Self-Consistency Wang et al. (2022b), and Graph-of-Thought Besta et al. (2024) enhance single-agent reasoning by decomposing tasks or broadening candidate exploration. Self-Refine Madaan et al. (2023) introduces iterative self-correction, and MemoryBank Zhong et al. (2024) augments planning with historical knowledge retrieval. However, these methods are designed for single-agent settings with full or near-complete observability. In multi-agent scenarios under partial observability, the core challenge shifts from individual reasoning quality to inter-agent coordination under fragmented information. That is a problem that single-agent planning enhancements do not address.
Extended Related Work on LLM-based Multi-Agent Systems.
The emergence of Large Language Models (LLMs) has revolutionized multi-agent collaboration. Early frameworks such as CAMEL Li et al. (2023a) and AutoGen Wu et al. (2023) demonstrated the potential of communicative agents to solve open-ended tasks through dynamic role-playing. Subsequent systems introduced structured enhancements to improve stability and capability. For instance, MetaGPT Hong et al. (2023) incorporates Standardized Operating Procedures (SOPs) to streamline workflows, while Voyager Wang et al. (2023) utilizes an iterative curriculum with persistent code libraries for embodied control. Despite their success in observable environments (e.g., software engineering, Minecraft), these frameworks typically operate under the assumption of global observability or unrestricted communication, rendering them less effective in spatially distributed, bandwidth-constrained real-world scenarios. Recent research targeting the information imbalance caused by spatial constraints generally falls into two categories:
  1. 1.
    Hierarchical aggregation: Works such as FinCon Yu et al. (2024), MLPO Estornell et al. (2025), and HALO Hou et al. (2025) employ hierarchical architectures. These systems designate specific agents (e.g., central hubs) to aggregate incomplete information, constructing a pseudo-global view to compensate for limited local perception. However, this centralization introduces critical bottlenecks: it relies heavily on central nodes and often incurs prohibitive communication overheads as the network scales.
  2. 2.
    Individual reasoning enhancement: A complementary line of research focuses on enhancing individual reasoning capabilities. Theory-of-Mind (ToM) approaches Li et al. (2023b) enable agents to infer the beliefs and intents of others, while frameworks like LAMEN Davidson et al. (2024) utilize structured reasoning chains to predict hidden states. While effective for small-scale interactions, these methods often struggle with the spatiotemporal uncertainty inherent in complex topologies, where agents must simultaneously reason about spatial neighbors and future dynamics under strict latency constraints.
  3. 3.
    Swarm-based optimization: Recent works apply particle swarm optimization to LLM systems. Model Swarms Feng et al. (2024) searches the model weight space over 50 iterations with per-particle validation-set evaluation to produce adapted expert checkpoints. SwarmAgentic Zhang et al. (2025) searches the system configuration space over 10 iterations, jointly optimizing agent functionalities and collaboration structures through failure-driven refinement. While both methods achieve dynamic configuration evolution, this adaptation occurs during an offline optimization loop that requires a globally accessible fitness function evaluated over the entire validation set.
In contrast, MACRO-LLM proposes a fully decentralized negotiation mechanism that leverages mean-field statistics and semantic strategies, effectively balancing local perception with global coordination goals.
Extended Related Work on MARL
To tackle spatial partial observability, the Decentralized Partially Observable Markov Decision Process (Dec-POMDP) serves as the foundational modeling framework Oliehoek et al. (2016). Under this model, agents must act based on local histories rather than global states. To mitigate the non-stationarity inherent in decentralized learning, the Centralized Training with Decentralized Execution (CTDE) paradigm has become standard. Recent works have further incorporated memory augmentation mechanisms Kia et al. (2024) and theoretical guarantees Bernstein et al. (2005) to approximate global state value functions. However, these methods typically require fitting a value function to a specific state space, limiting their transferability.
Scalability remains a critical bottleneck in MARL. Mean-field MARL offers a promising direction by approximating agent-to-agent interactions via the average effect of the neighborhood Yang et al. (2018). Subsequent studies have extended this to accommodate agent heterogeneity Ganapathi Subramanian et al. (2020); Gu et al. (2025) and robustness Wang et al. (2022a); Anand et al. (2024). Similarly, MACRO-LLM adopts the mean-field principle in its Negotiator module to compress spatial information. However, unlike MARL which learns a mean-field policy through extensive episodes, our approach computes mean-field statistical features explicitly as prompts for the LLM, enabling inference-time reasoning without gradient updates.
To improve sample efficiency, advanced policy optimization methods such as Multi-Agent PPO (MAPPO) Schulman et al. (2017) and Model-based Decentralized Policy Optimization (DMPO) Ma et al. (2024) have been developed. Furthermore, adaptive communication protocols like IC3Net Singh et al. (2018) introduce gating mechanisms to decide when to communicate. Despite these advances, the fundamental reliance on trial-and-error interaction makes MARL computationally intensive. As noted by Huh and Mohapatra (2023) and Oroojlooy and Hajinezhad (2023), MARL policies often struggle to generalize to unseen topologies or agent numbers, necessitating the retraining costs that our LLM-based framework aims to eliminate.
Appendix B Detailed Algorithm of Proposal Generation
Input: Observation $o_{n}^{t}$; strategies $\Pi_{n,t}^{\text{temp}}$ and $\Pi_{n,t}^{\text{spatial}}$; maximum attempts $Att_{\text{max}}$; rollout horizon $K$
Output: Verified proposal $P_{n}^{t}=(o_{n}^{t},{a_{n}^{t}},\{{a_{n\rightarrow m}^{t}}\}_{m\in\mathcal{N}_{n}})$
[…]
2 Initialize candidate action with highest reward score by $a_{n}^{t,0}\leftarrow\arg\max_{a}\mathsf{InferAction}(o_{n}^{t},\mathcal{A}_{n},\Pi_{n,t}^{\text{temp}},\Pi_{n,t}^{\text{spatial}})$ and ensure $R_{n}^{t}(o_{n}^{t},a_{n}^{t,0})\geq R_{n}^{t-1}(o_{n}^{t-1},a_{n}^{t-1})$;
3 while $att<Att_{max}$ do
4   Initialize rollout trajectory $\tau_{n}^{t,att}=(o_{n}^{t},a_{n}^{t,att},R_{n}^{t})$;
5 foreach $m\in\mathcal{N}_{n}$ do
6      Obtain observable agents’ candidate action ${a_{n\rightarrow m}^{t,att}}\leftarrow\mathsf{InferNeighborAction}(o_{n}^{t},{a_{n}^{t,att}},\Pi_{n,t}^{\text{spatial}})$;
7  Construct candidate proposal $P_{n}^{t,att}\leftarrow(o_{n}^{t},{a_{n}^{t,att}},\{{a_{n\rightarrow{m}}^{t,att}}\}_{m\in\mathcal{N}_{n}})$;
8 for Verified time step $k=1,\cdots,K$ in a rollout do
9      Predict next observation $o_{n}^{t+k,att}$ with candidate proposal $P_{n}^{t,att}$;
10    if $o_{n}^{t+k,att}$ violate constraints then
11         Append $P_{n}^{t,att}$ to Candidate_Proposals;
12       $att=att+1$;
13         Update candidate action $a_{n}^{t,att}$ slightly and ensure $R_{n}^{t}(o_{n}^{t},a_{n}^{t,0})\geq R_{n}^{t-1}(o_{n}^{t-1},a_{n}^{t-1})$;
14         Break;
15
16     Infer actions $a_{n\rightarrow m}^{t+k,att}$ on time $t+k$;
17      Add $(a_{n}^{t+k,att},o_{n}^{t+k,att},R_{n}^{t+k,att})$ to rollout trajectory $\tau_{n}^{t,att}$;
18
19 if $P_{n}^{t,{att}}$ has been verified over $k$ time steps then
20    return Optimal $P_{n}^{t}\leftarrow P_{n}^{t,att}$.
21
22Select the proposal with the highest cumulative reward in its rollout from Candidate_Proposals as $P_{n}^{t}$;
23 return $P_{n}^{t}$;
Algorithm 1 Pseudocode of proposal generation via rollout-simulated verification.
In this section, we detail the algorithmic procedure for the CoProposer module. The core mechanism, Proposal Generation via Rollout-Simulated Verification, is outlined in Algorithm 1.
This process aims to find an action sequence that maximizes cumulative reward while strictly adhering to safety constraints. To ensure computational efficiency, the algorithm employs an early exit mechanism: if a generated proposal successfully passes verification over the entire horizon $k$, it is immediately adopted as the proposal $P_{n}^{t}$, terminating the search loop.
The computational complexity of the CoProposer is dominated by the number of LLM inference calls. Let $Att_{\text{max}}$ be the maximum number of validation attempts and $K$ be the rollout horizon. In the worst-case scenario (where valid proposals are hard to find), the complexity is $\mathcal{O}(Att_{\text{max}}\cdot k)$. To mitigate latency in real-world applications, we set $Att_{\text{max}}=10$ for all tasks. The binary simplification of rewards beyond $t+1$ further reduces the overhead of reward calculation.
Appendix C Scope of Theoretical Assumptions
Eq. 4 analyzes the mean-field approximation error under the domain’s true reward function $R_{n}$, i.e., the error introduced by substituting the true neighborhood states $s^{t}_{\mathcal{N}_{n}}$ with the mean-field estimate $\mu_{n}$. We clarify the scope and assumptions below.
Domain Satisfaction.
In our two evaluation domains, the domain reward functions are governed by continuously differentiable physical dynamics: kinematic equations relating headway and velocity deviations in CPP, and epidemiological SIR-family dynamics governing infection rates in PC. Both operate over bounded state-action spaces with finite horizons, satisfying the smoothness requirement for the Taylor expansion argument. The constant $L_{2}$ is therefore determined by the continuity properties of these domain-specific dynamics.
Scope Clarification.
Eq. 4 characterizes the theoretical relationship between neighborhood variance $\sigma^{2}_{n}$ and the mean-field approximation quality under the domain’s true reward function. This analysis establishes $\sigma^{2}_{n}$ as a principled uncertainty indicator, motivating the framework’s design choice of transmitting $\sigma^{2}_{n}$ during negotiation. In practice, the LLM’s rollout-based reward estimation introduces a separate source of approximation that is outside the scope of Eq. 4. This estimation gap is mitigated through progressive constraint relaxation in rollout verification (Sec. 3.2), which simplifies reward signals beyond $t+1$ to binary feasibility indicators, reducing sensitivity to prediction imprecision.
Appendix D Proof of Theorem 1
Proof.
When the neighborhood $\mathcal{N}_{n}$ is separated into $P$ disjoint subgroups $\{\mathcal{N}_{n}^{(p)}\}_{p=1}^{P}$ and apply Eq. 4 independently within each subgroup. For each subgroup $p$, define the local weighted mean $\mu_{n}^{(p)}=\sum_{m\in\mathcal{N}_{n}^{(p)}}w_{n,m}^{(p)}\,s_{m}^{t}$ and local weighted variance $\sigma_{n}^{2,(p)}=\sum_{m\in\mathcal{N}_{n}^{(p)}}w_{n,m}^{(p)}\,(s_{m}^{t}-\mu_{n}^{(p)})^{2}$, where $w_{n,m}^{(p)}=w_{n,m}/W^{(p)}$ are the renormalized weights within subgroup $p$ and $W^{(p)}=\sum_{m\in\mathcal{N}_{n}^{(p)}}w_{n,m}$. By the same argument as in Eq. 4 (Taylor expansion with first-order cancellation within each subgroup), the contribution of subgroup $p$ to the approximation error is bounded by:
$\begin{split}|\Delta\mathcal{R}_{n}^{(p)}|&\leq\frac{L_{2}}{2}\sum_{m\in\mathcal{N}_{n}^{(p)}}(s_{m}^{t}-\mu_{n}^{(p)})^{2}\\ &\leq\frac{L_{2}}{2\,w_{\min}}\sum_{m\in\mathcal{N}_{n}^{(p)}}w_{n,m}\,(s_{m}^{t}-\mu_{n}^{(p)})^{2}.\end{split}$ (6)
Since the subgroups are disjoint, the total approximation error is bounded by:
$\begin{split}&\left|\mathcal{R}_{n}(\mathbf{s}_{\mathcal{N}_{n}}^{t})-\mathcal{R}_{n}(\boldsymbol{\mu}_{n}^{(1:P)})\right|\leq\sum_{p=1}^{P}|\Delta\mathcal{R}_{n}^{(p)}|\\ \leq&\frac{L_{2}}{2\,w_{\min}}\sum_{p=1}^{P}\sum_{m\in\mathcal{N}_{n}^{(p)}}w_{n,m}\,(s_{m}^{t}-\mu_{n}^{(p)})^{2}.\end{split}$ (7)
By the law of total variance Casella and Berger (2024), for any partition of a weighted dataset into $P$ groups:
$\begin{split}&\underbrace{\sum_{m\in\mathcal{N}_{n}}w_{n,m}\,(s_{m}^{t}-\mu_{n})^{2}}_{\sigma_{n}^{2}\text{ (total variance)}}\\ =&\underbrace{\sum_{p=1}^{P}W^{(p)}\sigma_{n}^{2,(p)}}_{\text{within-group variance}}+\underbrace{\sum_{p=1}^{P}W^{(p)}(\mu_{n}^{(p)}-\mu_{n})^{2}}_{\text{between-group variance}\;\geq\;0},\end{split}$ (8)
where $W^{(p)}=\sum_{m\in\mathcal{N}_{n}^{(p)}}w_{n,m}$. Since the between-group variance is non-negative:
$\sum_{p=1}^{P}W^{(p)}\sigma_{n}^{2,(p)}\leq\sigma_{n}^{2}.$ (9)
Furthermore, noting that $\sum_{m\in\mathcal{N}_{n}^{(p)}}w_{n,m}\,(s_{m}^{t}-\mu_{n}^{(p)})^{2}=W^{(p)}\sigma_{n}^{2,(p)}$, we can obtain:
$\begin{split}&\sum_{p=1}^{P}\sum_{m\in\mathcal{N}_{n}^{(p)}}w_{n,m}\,(s_{m}^{t}-\mu_{n}^{(p)})^{2}\\ =&\sum_{p=1}^{P}W^{(p)}\sigma_{n}^{2,(p)}\leq\sigma_{n}^{2}.\end{split}$ (10)
Substituting back into (7):
$\begin{split}&\left|\mathcal{R}_{n}(\mathbf{s}_{\mathcal{N}_{n}}^{t})-\mathcal{R}_{n}(\boldsymbol{\mu}_{n}^{(1:P)})\right|\\ \leq&\frac{L_{2}}{2\,w_{\min}}\sum_{p=1}^{P}W^{(p)}\sigma_{n}^{2,(p)}\leq\frac{L_{2}}{2\,w_{\min}}\cdot\sigma_{n}^{2}\\ =&\frac{L_{2}}{2\,w_{\min}}\cdot|\mathcal{N}_{n}|\cdot\bar{\sigma}_{n}^{2},\end{split}$ (11)
where $\bar{\sigma}_{n}^{2}=\sigma_{n}^{2}/|\mathcal{N}_{n}|$ is the per-neighbor average variance. The inequality is strict whenever the subgroup means $\mu_{n}^{(p)}$ differ, since the between-group variance $\sum_{p}W^{(p)}(\mu_{n}^{(p)}-\mu_{n})^{2}>0$ in that case. Therefore, if the topology can be divided into several disjoint group, the error bound is tighter. ∎
Appendix E Proof of Theorem 2 (Properties of the Negotiation Loop)
Proof.
$\mathcal{G}=(\mathcal{V},\mathcal{E})$ represents a connected communication graph, and $r$ denotes the maximum number of negotiation rounds.
Bounded Termination.
At each round $\ell$, the negotiation loop evaluates one of two termination conditions:
  • Early stopping via local consensus: If $\|P_{n}^{(\ell)}-P_{m}^{(\ell)}\|<\delta$ for all $m\in\mathcal{N}_{n}$, where $\delta$ is a predefined task-specific threshold (e.g., $\delta=0.1\,\text{m/s}^{2}$ for CPP), the agent terminates negotiation.
  • Maximum round limit: The loop terminates after at most $r$ rounds regardless of consensus status.
Since both conditions are finitely bounded, termination is guaranteed for every agent $n\in\mathcal{V}$.
Monotonic Non-Decrease of Proposal Quality.
At the beginning of each negotiation round $\ell+1$, the best proposal from the previous round, $P_{n}^{(\ell)}$, is retained in the candidate set for the current round. That is, the candidate set at round $\ell+1$ is:
$\mathcal{C}_{n}^{(\ell+1)}=\{P_{n}^{(\ell)}\}\cup\mathcal{C}_{n}^{\text{new}},$ (12)
where $\mathcal{C}_{n}^{\text{new}}$ denotes the set of newly generated proposals incorporating neighbor information received at round $\ell+1$.
Since $P_{n}^{(\ell)}\in\mathcal{C}_{n}^{(\ell+1)}$ by the Design Invariant, and agent $n$ selects $P_{n}^{(\ell+1)}=\arg\max_{P\in\mathcal{C}_{n}^{(\ell+1)}}\mathcal{R}(\tau_{n}(P))$, we have $\mathcal{R}(\tau_{n}^{(\ell+1)})=\max_{P\in\mathcal{C}_{n}^{(\ell+1)}}\mathcal{R}(\tau_{n}(P))\geq\mathcal{R}(\tau_{n}(P_{n}^{(\ell)}))\mathcal{R}(\tau_{n}^{(\ell)}).$ Therefore, the sequence $\{\mathcal{R}(\tau_{n}^{(\ell)})\}_{\ell=0}^{r}$ is monotonically non-decreasing.
The cumulative rollout reward is upper-bounded by the maximum achievable return over the rollout horizon $k$:
$\mathcal{R}(\tau_{n}^{(\ell)})\leq\sum_{\iota=0}^{k}\gamma^{\iota}R_{\max}=R_{\max}\cdot\frac{1-\gamma^{k+1}}{1-\gamma},$ (13)
where $R_{\max}=\sup_{o_{n},a_{n}}|R_{n}(o_{n},a_{n})|$ is the single-step reward bound. Therefore, $\{\mathcal{R}(\tau_{n}^{(\ell)})\}_{\ell=0}^{r}$ is a bounded, monotonically non-decreasing sequence of real numbers, which necessarily converges. ∎
Appendix F Equations for Welford-based State Aggregation
The update rules for the weighted mean and variance are defined as follows:
$\mu_{n}^{\prime}=\mu_{m}+\frac{w_{n}}{W_{n}}(s_{n}^{t}-\mu_{m}),$ (14)
${\sigma_{n}^{2}}^{\prime}=\frac{W_{m}}{W_{n}}\sigma_{m}^{2}+\frac{w_{n}}{W_{n}}(s_{n}^{t}-\mu_{m})(s_{n}^{t}-\mu_{n}^{\prime}),$ (15)
where $W_{n}=W_{m}+w_{n}$ represents the updated weight. The aggregated mean $\mu_{n}^{\prime}$ and variance $\sigma_{n}^{2}$$^{\prime}$ are then abstracted into a textual description of a virtual node within the query prompt, enabling the LLM to analyze trends in the unobserved states. While equations (14) and (15) illustrate the atomic update, in practice, agents apply this iteratively or employ parallel variance merger algorithms Chan et al. (1983) when aggregating disjoint sets.
Appendix G Simulation Environments and Task Configurations
Table 3 summarizes the key settings of the MACRO-LLM framework used across all experiments.
Table 3: Framework hyperparameters.
Parameter Value
Rollout horizon $k$ 2
Max negotiation rounds $r$ 3
Max proposal attempts $\text{Att}_{\max}$ 10
Mean-field weight $w_{n,m}$ Inverse hop distance
Consensus threshold $\epsilon$ 0.02 (CPP) / 0 (PC)
Temperature / top-p 0.3 / 1.0
In MACRO-LLM, the rollout horizon $k=2$ balances verification depth against prompt length: the first rollout step performs precise constraint verification on the immediate next state, while the second step applies relaxed feasibility checks to ensure short-term safety (Sec. 3.3). The negotiation rounds $r=3$ are set based on network diameter $D$: since each agent begins with 1-hop neighbor information and each round propagates aggregated statistics by one additional hop, $r$ rounds yield effective coverage of $r+1$ hops. For the PC topologies ($D\leq 4$), $r=3$ provides full network coverage. For the CPP linear topology ($D=7$), the topology-aware partitioning into predecessor and follower subgroups (Theorem 1) reduces the effective propagation distance to $D/2$ per direction, which $r=3$ sufficiently covers.
We evaluate our framework on two complementary domains, CPP and PC, to assess its collaborative performance and generalization capabilities, utilizing high-fidelity simulators for each task. The evaluation covers continuous and discrete actions, linear and graph topologies, fine-grained and long-horizon planning.
G.1 Cooperative Platoon Planning (CPP)
The CPP task is simulated using the Simulation of Urban Mobility (SUMO) [^1] SUMO: https://eclipse.dev/sumo/, licensed under EPL 2.0. (Version 1.21.0).
  • Topology: We model a platoon of eight vehicles ($N=8$) on a single-lane highway segment. The simulation runs for a duration of 60s with 120 time steps in total. Each agent operates with limited perception, accessing only the state information (e.g., headway, velocity) of its immediately adjacent vehicles (predecessor and follower). Due to the linear topology, each agent’s neighborhood naturally partitions into two disjoint subgroups—predecessor and follower—with separate mean-field statistics computed for each direction (Theorem 1).
  • Objective: The control objective is for agents to autonomously adjust acceleration to stabilize the platoon, converging rapidly to an optimal headway of 20 m and a target velocity of 15 m/s. The consensus threshold is set to $\epsilon=0.02$ for the CPP task.
  • Scenarios: We evaluate performance in two dynamic settings: (1) Scenario 1: Catch-up: The leading vehicle accelerates to close the gap with a target vehicle, testing the platoon’s ability to re-establish stable formation under high-speed conditions. (2) Scenario 2: Slow-down: The leading vehicle decelerates from a high initial speed, challenging the followers to execute coordinated braking without collisions or oscillation.
G.2 Pandemic Coordination (PC)
Figure 6: Communication topology of cities for pandemic networks. (a) Helsinki (b) Hong Kong (c) New York.
The PC task is simulated using the open-source Pandemic Simulator [^2] Pandemic Simulator: https://github.com/SonyResearch/PandemicSimulator, licensed under Apache-2.0. Kompella* et al. (2020).
  • Topology: We model the urban environment as a graph of seven agents representing distinct facility types: Home, Office, School, Hospital, Retail, Restaurant, and Government. To construct realistic urban topologies, we utilize Google Maps to identify clusters of functionally similar buildings within specific urban blocks, marking their centroids as individual agents. Communication links are established based on spatial proximity, forming a connected topology as shown in Fig. 6. Since the urban graph does not admit a natural disjoint partition of each agent’s neighborhood, mean-field statistics are computed over the full neighborhood without subgroup decomposition. The simulation spans 120 days, during which agents operate under partial observability regarding global infection rates.
  • Objective: The primary objective is for agents to dynamically adjust regulatory policies (e.g., gathering restrictions, business hours) to minimize the spread of the virus while maintaining essential socio-economic activities. The consensus threshold is set to $\delta=0$ for the PC task.
  • Scenarios: We evaluate the framework’s generalization capabilities across three distinct urban settings with varying population scales: (1) Helsinki: A standard community setting with a population of 500. (2) Hong Kong: A denser urban setting with a population of 1000. (3) New York: A high-density complex topology with a population of 1500.
Appendix H Baseline Methods
To evaluate MACRO-LLM, we implement and adapt the following baselines for both the CPP and PC tasks.
H.1 MARL Baselines
We compare against traditional Multi-Agent Reinforcement Learning (MARL) approaches: (1) DPPO Ma et al. (2024): Multi-agent Decentralized Proximal Policy Optimization; (2) DMPO Ma et al. (2024): Multi-agent Model-based Decentralized Policy Optimization; (3)IC3Net Singh et al. (2018): A classic approach for learning when to communicate in multi-agent environments.
H.2 LLM-based MAS Baselines
We compare against recent frameworks utilizing LLMs for collaboration: (1) ToM-Belief Li et al. (2023b): A MAS that employs second-order Theory of Mind for introspection via belief state updates and environmental feedback; (2) ChatEval Chan et al. (2023): Adopts the Simultaneous-Talk-with-Summarizer protocol (Algorithm 3), which is structurally similar to our task setting; (3) LAMEN Davidson et al. (2024): A framework that constructs internal “mental notes” alongside public messages to facilitate strategic negotiation.
H.3 Adaptation, Fairness, and System Design
To ensure a fair and reproducible comparison, we maintain experimental consistency across all LLM-based methods. In MACRO-LLM, the LLM performs all strategic reasoning: generating collaborative proposals, predicting future observations during rollout verification, evaluating proposal conflicts and updating spatial strategies in the Negotiator, and diagnosing performance drops in the Introspector. Deterministic computations independent of the LLM reasoning process, such as mean-field statistics (Eqs. 23) and drift intensity (Eq. 5), are offloaded to an external numerical module to ensure precision. All inference-dependent computations, including state prediction and constraint evaluation during rollout verification, are performed entirely by the LLM. LLM outputs are validated against a structured XML schema; outputs failing validation trigger an automatic retry (up to 3 attempts).
All LLM-based baselines (ToM-Belief, ChatEval, LAMEN) share the same backbone model (GPT-4o), environmental interfaces, external numerical module, schema validation, and retry mechanism as MACRO-LLM. Methods differ only in their prompting strategies and inter-agent communication protocols (e.g., mental notes generation in LAMEN, belief updates in ToM-Belief). This ensures that performance differences reflect methodological contributions rather than implementation advantages.
Appendix I Case Study: Hierarchical Deployment for Safety-Critical Domains
As a general-purpose strategic coordination framework, MACRO-LLM outputs high-level collaborative directives rather than low-level control commands. For domains with real-time safety constraints, this property allows MACRO-LLM to naturally integrate into a hierarchical deployment paradigm Liu et al. (2025); Wen et al. (2024), where the strategic layer and execution layer operate at different temporal granularities. In this section, we describe the deployment architecture and experimentally validate its feasibility through a decision frequency analysis on the Cooperative Platoon Planning task.
Figure 7: Hierarchical deployment paradigm for MACRO-LLM. The strategic coordination layer executes a full negotiation cycle every $\Delta$ steps to produce collaborative directives, while the domain-specific execution layer operates at every step, applying the cached strategic plan when the state is safe or triggering a safety controller override when a threshold is breached.
I.1 Hierarchical Deployment Paradigm
As illustrated in Fig. 7, the hierarchical paradigm consists of two layers:
  1. 1.
    Strategic Coordination Layer (Second-level granularity). MACRO-LLM operates at a coarser temporal granularity, executing a full negotiation cycle among agents every $\Delta$ time steps. At each decision epoch, it produces strategic directives by reasoning about unobservable states and negotiating with neighboring agents. Between decision epochs, the execution layer carries out the planned actions.
  2. 2.
    Domain-Specific Execution Layer (Millisecond-level granularity). A lightweight controller operates at every time step, performing a safety threshold check on the current state. If the state remains within the safe region, the controller executes the cached strategic plan from the most recent MACRO-LLM decision epoch. If a safety threshold is breached, a domain-specific controller (e.g., PID for CPP) immediately overrides the planned action to ensure system safety.
MACRO-LLM requires no architectural modification to operate within this paradigm—its outputs are inherently strategic directives that can be consumed by any domain-specific execution layer.
I.2 Decision Frequency Analysis
To validate the feasibility of hierarchical deployment, we evaluate MACRO-LLM on the CPP Catch-up scenario under varying decision intervals $\Delta\in\{1,3,5,10\}$, where $\Delta=1$ corresponds to the original setting used in our main experiments. At each decision epoch, MACRO-LLM produces a strategic action spanning the subsequent $\Delta$ time steps. Between decision epochs, the execution layer applies the most recently planned action while monitoring safety constraints via a proportional headway-tracking controller with emergency override. Table 4 presents the results alongside representative baselines from Table 1 for reference.
Table 4: Decision frequency analysis on the CPP Catch-up scenario. DPPO and LAMEN from Table 1 for reference. All metrics $\downarrow$.
Config. Red. RMSE-H RMSE-V SD-H SD-V
DPPO 1.603 0.978 0.994 0.474
LAMEN 1.891 0.648 3.200 0.601
Hier. $\Delta=10$ 90% 1.443 0.623 0.634 0.225
Hier. $\Delta=5$ 80% 1.319 0.721 0.524 0.281
Hier. $\Delta=3$ 67% 1.408 0.596 0.564 0.234
MACRO $\Delta=1$ 0% 1.237 0.721 0.409 0.203
Hier.: hierarchical deployment;
$\Delta$: decision interval (time steps between MACRO-LLM invocations)
Red.: LLM calls reduced relative to $\Delta=1$.
All hierarchical configurations maintain safe operation without any collisions, confirming that the execution layer’s safety override mechanism provides reliable protection between decision epochs. As shown in Table 4, even with 90% fewer LLM invocations ($\Delta=10$), the hierarchical deployment consistently outperforms both DPPO and LAMEN across all metrics. The SD-H of hierarchical deployment ($\Delta=10$) is 36.2% lower than that of DPPO and 80.2% lower than that of LAMEN. This demonstrates that MACRO-LLM’s coordination reasoning provides substantial value even at highly reduced decision frequencies, as the strategic directives continue to guide inter-agent collaboration during the intervals between LLM invocations. Furthermore, increasing the decision frequency yields progressive improvements in coordination quality: SD-H improves from 0.634 ($\Delta=10$) to 0.409 ($\Delta=1$), a 35.5% reduction, confirming that the performance gains stem from MACRO-LLM’s collaborative reasoning. These results indicate that MACRO-LLM can be effectively deployed within a hierarchical paradigm, achieving competitive coordination quality while substantially reducing computational overhead.
Appendix J Detailed Evaluation Metrics
We provide the detailed definitions and mathematical formulations of the metrics used in our experiments.
J.1 CPP Task Metrics
Following previous works Desjardins and Chaib-draa (2011); Van Arem et al. (2006), we evaluate the platoon efficiency and stability using:
  • Root Mean Square Error (RMSE-H/V): Measures the deviation between the actual headway/velocity and the target values, averaged across all following vehicles ($i=1\dots N$) over the simulation steps ($t=1\dots T$). For example, RMSE of Headway (RMSE-H) is calculated as:
    $\text{RMSE-H}=\sqrt{\frac{1}{N\cdot T}\sum_{t=1}^{T}\sum_{i=1}^{N}(h_{i}^{t}-h_{\text{target}})^{2}},$ (16)
    where $h_{i}^{t}$ is the headway of vehicle $i$ at time $t$, and $h_{\text{target}}$ is the optimal safety distance. RMSE-V is computed similarly using velocity deviations as follows:
    $\text{RMSE-V}=\sqrt{\frac{1}{N\cdot T}\sum_{t=1}^{T}\sum_{i=1}^{N}(v_{i}^{t}-v_{\text{leader}})^{2}},$ (17)
    where $v_{\text{leader}}$ denotes the velocity of the leading vehicle.
  • Standard Deviation (SD-H/V): Measures the collaborative consistency of the platoon across all agents. Specifically, at each time step $t$, we calculate the standard deviation of the headway/velocity across all $N$ vehicles. The final reported metric is the temporal average of these spatial standard deviations over the duration $T$:
    $\text{SD-H}=\frac{1}{T}\sum_{t=1}^{T}\sqrt{\frac{1}{N}\sum_{i=1}^{N}(h_{i}^{t}-\bar{h}^{t})^{2}},$ (18)
    where $\bar{h}^{t}=\frac{1}{N}\sum_{j=1}^{N}h_{j}^{t}$ is the average headway of the entire platoon at time $t$. A lower SD value indicates that all agents are maintaining consistent states simultaneously, reflecting high coordination stability. SD-V is computed similarly using velocity deviations as follows:
    $\text{SD-V}=\frac{1}{T}\sum_{t=1}^{T}\sqrt{\frac{1}{N}\sum_{i=1}^{N}(v_{i}^{t}-\bar{v}^{t})^{2}}.$ (19)
J.2 PC Task Metrics
Following Kompella et al. Kompella* et al. (2020), we employ the following metrics to assess containment efficacy. Let $P_{total}$ be the total population, and $I^{t},C^{t},D^{t}$ be the count of infected, critical, and dead cases on day $t$, respectively.
  • Normalized Infection (I_n): The ratio of the cumulative number of infected cases to the total population:
    $I\_n=\frac{\sum_{t}\Delta I_{new}^{t}}{P_{total}}.$ (20)
  • Normalized Peak Infection (PI_n): Measures the highest pressure on the healthcare system:
    $PI\_n=\frac{\max_{t}(I^{t})}{P_{total}}.$ (21)
  • Normalized Dead (D_n): The ratio of total dead cases to the total population at the end of the episode ($t_{end}$):
    $D\_n=\frac{D^{t_{end}}}{P_{total}}.$ (22)
  • Pandemic Duration (PD): The time span (in days) from the first confirmed infection ($t_{start}$) until the pandemic is effectively contained:
    $PD=t_{end}-t_{start},$ (23)
    where $t_{end}=\min\{t\mid I^{t}+C^{t}=0,\forall\tau>t\}$.
Appendix K Analysis of Model Generalizability
Table 5: Generalizability analysis of MACRO-LLM across different LLM backbones. S1/S2: Catch-up/Slow-down scenarios of CPP tasks. C1-C3: PC topologies for Helsinki, Hong Kong, and New York.
Metrics ($\downarrow$)
Tasks Base Models RMSE-H RMSE-V SD-H SD-V
CPP S1 DeepSeek-V3.1 1.528 0.553 0.611 0.190
Llama-3.3-70B 1.412 0.779 0.645 0.365
Mistral-3.2-24B 1.891 0.784 0.997 0.326
Qwen3-flash 3.521 1.910 2.935 0.579
GPT-4o-mini 2.906 1.667 1.858 0.936
GPT-4o 1.237 0.721 0.409 0.203
S2 DeepSeek-V3.1 0.649 3.388 0.274 0.417
Llama-3.3-70B 1.163 3.736 0.889 1.224
Mistral-3.2-24B 1.842 3.523 1.157 0.916
Qwen3-flash 3.871 3.678 2.946 0.974
GPT-4o-mini 2.352 3.999 1.598 1.171
GPT-4o 0.496 3.310 0.128 0.238
Tasks Base Models I_n PI_n D_n PD
PC C1 DeepSeek-V3.1 0.098 0.038 0.002 46
Llama-3.3-70B 0.013 0.008 0.000 14
Mistral-3.2-24B 0.043 0.014 0.003 31
Qwen3-flash 0.010 0.008 0.002 16
GPT-4o-mini 0.010 0.010 0.000 17
GPT-4o 0.011 0.008 0.000 10
C2 DeepSeek-V3.1 0.008 0.005 0.001 22
Llama-3.3-70B 0.012 0.005 0.000 16
Mistral-3.2-24B 0.009 0.004 0.001 22
Qwen3-flash 0.008 0.006 0.001 18
GPT-4o-mini 0.006 0.004 0.000 24
GPT-4o 0.008 0.004 0.001 23
C3 DeepSeek-V3.1 0.003 0.003 0.000 11
Llama-3.3-70B 0.013 0.006 0.000 25
Mistral-3.2-24B 0.009 0.006 0.000 14
Qwen3-flash 0.003 0.003 0.000 13
GPT-4o-mini 0.003 0.003 0.000 13
GPT-4o 0.004 0.003 0.000 25
We investigate the robustness of MACRO-LLM by instantiating it with diverse LLM backbones, including DeepSeek-V3.1 Liu et al. (2024), Llama-3.3-70B-Instruct Grattafiori et al. (2024), Mistral-Small-3.2-24B-Instruct, Qwen3-flash Bai et al. (2023), GPT-4o-mini, and GPT-4o. The results are detailed in Table 5. In the CPP task, performance gaps across backbones are primarily attributable to the fine-grained continuous control demands: GPT-4o and DeepSeek-V3.1 reduce RMSE-H by over 47.4% compared to GPT-4o-mini in the Catch-up scenario. Conversely, in the PC task, Qwen3-flash and GPT-4o-mini achieve containment efficacy comparable to GPT-4o across all three topologies, as this task relies more on strategic planning than on numerical precision. This contrast confirms that the observed performance variation stems from task-specific precision requirements rather than framework limitations, and that MACRO-LLM’s structured negotiation mechanism enables even lightweight models to accomplish complex multi-agent coordination tasks.
Appendix L Computational and Communication Scalability Analysis
We provide a theoretical and experimental analysis of MACRO-LLM, focusing on communication complexity and computational workload.
L.1 Communication Complexity Analysis
Following the definition by Yao Yao (1979), communication complexity is quantified by the total number of bits exchanged during the protocol. In our decentralized framework, we focus on the per-agent communication load. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be the agent communication graph. We denote the average node degree as $\bar{d}$ and the maximum negotiation rounds as $r$. The topology of the CPP task is linear ($\bar{d}\approx 2$). In the PC task, the topology represents an urban graph with average degree $\bar{d}_{\text{PC}}$. We assume sparse topologies where $\bar{d}$ does not grow linearly with the total number of agents $N$ (i.e., $\bar{d}\ll N$). As the network scales up, MACRO-LLM employs mean-field approximation to compress the states of unobservable agents into a single virtual node representation. Neglecting inference stochasticity, the per-agent message size is theoretically invariant to the network scale, which is represented as $M$. Consequently, the per-agent communication complexity can be represented as $O(\bar{d}\cdot r\cdot M)$. Since $\bar{d}$, $r$, and $M$ are independent of network scale $N$, the per-agent load remains constant as the network scales in theory. In other word, the per-agent communication load does not increase with the number of agents as long as the topology’s connectivity degree remains constant. The scalability results in Table 6 demonstrate as the platoon scales from $N=8$ to $N=32$, coordination metrics remain stable with no monotonic degradation.
L.2 Computational Efficiency and Latency
Given that the input prompt length remains bounded as analyzed above, the upper bound of the inference workload for a single agent does not increase with the number of agents. Under API-based deployment, the per-step latency is dominated by network round-trip overhead of remote API calls rather than the framework’s reasoning complexity itself. The per-step latency is dominated by API network overhead of remote LLM calls ($\sim$5.99s per call for GPT-4o). This overhead can be substantially mitigated through two orthogonal and composable strategies:
(1) Local deployment of lightweight models. Table 5 demonstrates that lightweight models such as GPT-4o-mini and Qwen3-flash achieve competitive coordination performance across both tasks. Locally deploying these models eliminates API network latency entirely, reducing per-call inference time by an order of magnitude compared to remote API access.
(2) Hierarchical deployment with reduced decision frequency. As demonstrated in Appendix I, MACRO-LLM maintains superior performance over all baselines even when decision frequency is reduced by 90% (Table 4), confirming that the framework’s coordination value does not depend on high-frequency invocation.
These two strategies are composable: combining local lightweight models with increased decision intervals can reduce the effective latency by over an order of magnitude relative to the default API-based configuration.
Furthermore, MACRO-LLM exhibits instant zero-shot adaptation to new scenarios, whereas MARL baselines necessitate extensive retraining when transferring to unseen topologies. For dynamic, open-world deployments where frequent retraining is impractical, this zero-shot capability represents a significant practical advantage that offsets the per-step inference cost.
L.3 Scalability Metrics Across Fleet Sizes
Fig. 3 represents the coordination performance for the CPP Catch-up scenario as the fleet size increases from $N=8$ to $N=32$. Table 6 reports the detailed metrics.
Table 6: Scalability analysis on CPP Catch-up with GPT-4o.
[…]
8 1.237 0.721 0.409 0.203
16 1.003 0.749 0.493 0.355
24 0.908 0.664 0.520 0.382
32 0.728 0.628 0.433 0.397
All metrics remain stable as $N$ increase to 32, with no monotonic degradation. The results indicate that the mean-field aggregation mechanism effectively compresses large-scale interactions without sacrificing coordination precision.
Appendix M Extended Ablation Analysis
In this section, we provide a detailed quantitative analysis of the ablation study.
M.1 Component Analysis in CPP Task
The removal of specific modules leads to critical performance degradation in the high-precision and highly dynamic CPP task:
[…]
M.2 Component Analysis in PC Task
A similar pattern holds for the PC task:
[…]
  • Impact of Negotiator: Excluding the Negotiator prolongs the pandemic to 32 days, demonstrating that the absence of consensus mechanisms leads to inconsistent and ineffective containment policies.
  • Impact of Introspector: Ablating the Introspector results in a pandemic duration approximately 44.4% longer than MACRO-LLM. This illustrates that agents without self-reflection are slow to adapt their strategies to the evolving infection dynamics.
Appendix N Analysis of Robustness
Table 7: Robustness analysis of MACRO-LLM with GPT-4o on the CPP task (5 independent runs). Results are reported as mean $\pm$ standard deviation.
Scenario RMSE-H ($\downarrow$) RMSE-V ($\downarrow$) SD-H ($\downarrow$) SD-V ($\downarrow$)
Catch-up 1.237 $\pm$ 0.068 0.721 $\pm$ 0.124 0.409 $\pm$ 0.034 0.203 $\pm$ 0.054
Slow-down 0.496 $\pm$ 0.044 3.310 $\pm$ 0.010 0.128 $\pm$ 0.054 0.238 $\pm$ 0.073
Table 8: Robustness analysis of MACRO-LLM with GPT-4o on the PC task (5 independent runs). Results are reported as mean $\pm$ standard deviation.
City I_n ($\downarrow$) PI_n ($\downarrow$) D_n ($\downarrow$) PD_n* ($\downarrow$)
Helsinki 0.011 $\pm$ 0.001 0.008 $\pm$ 0.000 0.000 $\pm$ 0.000 0.081 $\pm$ 0.024
Hong Kong 0.008 $\pm$ 0.004 0.004 $\pm$ 0.001 0.001 $\pm$ 0.001 0.190 $\pm$ 0.063
New York 0.004 $\pm$ 0.001 0.003 $\pm$ 0.000 0.000 $\pm$ 0.000 0.204 $\pm$ 0.117
* PD_n = PD / total simulation days, representing the proportion of pandemic duration relative
to the total evaluation period.
To evaluate the stability of MACRO-LLM, we conduct 5 independent runs with GPT-4o and report the mean and standard deviation in Tables 7 and 8. For the CPP task, MACRO-LLM exhibits low variance across all metrics. In the Catch-up scenario, the SD-H of $0.409\pm 0.034$ is 58.9% lower than that of DPPO, and even the upper bound ($0.443$) remains 55.4% lower, confirming that the stability advantage in Table 1 is not attributable to favorable randomness. Similarly, in the Slow-down scenario, the RMSE-H ($0.496\pm 0.044$) is 46.3% lower than that of LAMEN and 71.2% lower than that of DPPO across all runs. For the PC task, infection-related metrics (I_n, PI_n, D_n) remain tightly concentrated across all three topologies, with standard deviations at or near zero. While PD_n shows relatively higher variance in more complex topologies (e.g., New York: $0.204\pm 0.117$), this is expected given the larger state space and stochastic disease dynamics inherent to denser urban networks. These results confirm that MACRO-LLM achieves consistently superior performance with reliable reproducibility across independent trials.
Appendix O Analysis of Failure Cases and Solutions
To provide insights into the limitations of current LLMs in multi-agent control, we categorize the primary failure cases observed during our experiments and detail the corresponding mitigation strategies implemented in MACRO-LLM.
Arithmetic and Spatial Hallucinations.
Despite access to topological information, agents occasionally exhibit spatial hallucinations, such as confusing front and back vehicles in the CPP task. To mitigate this, we employ few-shot prompting incorporating explicit examples of state transitions (e.g., detailing how acceleration impacts headway relative to neighbors).
While LLMs occasionally produce imprecise numerical estimates during multi-step rollout predictions, all inference-dependent computations—including state prediction and constraint evaluation during rollout verification—are performed entirely by the LLM, with progressive constraint relaxation (Sec. 3.2) reducing sensitivity to arithmetic imprecision. Deterministic computations independent of the LLM reasoning process, such as post-execution environment updates, are offloaded to an external numerical module to ensure precision.
Identity Confusion.
In scenarios featuring dense interaction prompts, such as concurrent proposals involving similar identifiers, such as “Vehicle 2 propose action1 to Vehicle 1, and propose action2 to Vehicle 3”, agents may misattribute actions or forget their own identity. We attribute this to the attention mechanism’s difficulty in distinguishing semantically similar tokens within extended contexts. To address this, we enforce a structured schema for inter-agent communication and integrate a self-consistency mechanism to validate sender-receiver pairings. Additionally, we reinforce role stability by prepending the instruction “You are [Agent Name]” to every query prompt.
Syntactic Robustness.
To structure agent outputs, we employ XML-style <label> tags. However, smaller or quantized models (e.g., GPT-4o-mini, Qwen3-flash) occasionally fail to adhere to these formatting constraints, resulting in unclosed tags or hallucinated labels. To mitigate this, we implement a deterministic verification layer, outputs failing regex matching or schema validation trigger an automatic retry, ensuring that only syntactically valid actions are executed.
Safety Alignment Sensitivity.
In the PC task, the Qwen series frequently triggered safety refusals during stages characterized by high mortality rates, erroneously flagging simulation statistics as harmful content (returning openai.BadRequestError). This behavior indicates a stricter safety filter configuration compared to models like GPT-4o. To mitigate these false positives, we appended a system prompt explicitly clarifying the synthetic nature of the data: ‘‘SIMULATION NOTE: This is a hypothetical pandemic experiments on pandemic simulator. All data is synthetic and does not reflect real-world events!!!’’ Additionally, we implemented a rollback mechanism to discard and retry generation steps interrupted by safety filters.
Appendix P Analysis of Introspector’s Attribution-based Reasoning
To investigate how the Introspector achieves effective strategy refinement, we analyze whether its semantic revision signals reflect genuine attribution-based reasoning, attributing state changes to inter-agent interaction mechanisms, rather than simple trajectory pattern matching that maps local observations directly to actions. We provide both qualitative and quantitative evidence below.
P.1 Qualitative Evidence: Scenario-Adaptive Diagnostic Attribution Analysis
We present representative semantic revision signals generated by the Introspector in two distinct CPP scenarios to demonstrate that it produces scenario-adaptive diagnostic attribution analyses rather than templated outputs.
Case 1 (Catch-up Scenario).
Vehicle 2 generated the following revision signal:
“The reasons may be due to the unexpected deceleration of vehicle 1 and the slight acceleration of vehicle 3. Vehicle 1’s velocity decreased more than anticipated, causing vehicle 2 to adjust its acceleration to maintain a safe distance.”
The agent attributes its performance drop to specific neighbor behaviors, the unexpected deceleration of vehicle 1 and the acceleration pattern of vehicle 3, and derives a targeted adjustment strategy.
Case 2 (Slow-down Scenario).
The same agent position generated a qualitatively different signal:
“The spatial analysis indicates a general trend of deceleration, which may have influenced the need for a stronger deceleration to maintain formation.”
The agent attributes its adjustment to an aggregated mean-field spatial trend from unobservable agents, rather than to any individual neighbor’s behavior.
These two cases demonstrate fundamentally different attribution directions: Case 1 performs local neighbor-level attribution, while Case 2 reasons about collective spatial dynamics. This confirms that the Introspector adapts its diagnostic attribution analysis to the scenario’s underlying structure rather than applying a fixed reasoning pattern.
P.2 Quantitative Evidence: Diagnostic Attribution Statistics
To systematically evaluate whether the Introspector’s reflections constitute attribution reasoning, we analyzed three complete collaboration logs comprising all reflection instances across full simulation episodes. We clarify that the attribution explicitly references other agents’ behaviors, neighborhood statistical features, or inter-agent interaction effects as contributing factors to the observed performance change. Results show that 89.3% of reflections contain explicit attribution to other agents’ behaviors or neighborhood features, rather than self-trajectory patterns. This indicates that the Introspector systematically reasons about multi-agent interaction mechanisms rather than overfitting to local observation sequences.
P.3 Behavioral Coherence Verification
We further verify that the Introspector’s attribution reasoning translates into coherent behavioral adjustments. In the Catch-up case (Case 1 above), vehicle 2 identified that vehicle 1 decelerated unexpectedly (acceleration shifting from +1.8+1.8 to 1.3-1.3 m/s2) while vehicle 3 maintained low acceleration. The diagnostic attribution concluded that stabilization was needed to prevent headway violation. Consequently, vehicle 2 adjusted its own acceleration from 1.05-1.05 to 0.82-0.82 m/s2, a conservative correction aligned with the diagnosed cause, reducing deceleration intensity to avoid over-correction given vehicle 1’s unexpected behavior.
This behavioral coherence, where the direction and magnitude of action adjustments logically follow from the attribution diagnosis, is further corroborated by the ablation results in Figs. 4 and 5, where removing the Introspector leads to persistent oscillations in CPP ($1.9\times$ degradation in RMSE-H) and 44.4% longer pandemic duration in PC, confirming that attribution adaptation is essential for long-horizon coordination.
Appendix Q Core Prompts of MACRO-LLM
Q.1 CoProposer
Fig. 8 provides the core prompts of proposal generation with a rollout-simulated verification in CoProposer.
Refer to caption CoProposer: Proposal Generation Query You are [agent_name] in a [task_name]. This is negotiation round 0 with your neighbors. Your task is to [task objective]. Decision-Making Steps (Follow strictly, step by step):
Step 0: Initial Decision for [agent_name] at the 1st time step, and define urgency of the action: Normal, Warning, Urgent.
Step 1: Collaborative acceleration for [observable_agents] to achieve objective with [Temporal Strategy], [Spatial Strategy] and [Spatial Features].
Step 2-4: Validation by 2 Time-Step Simulation: (time step 2 and time step 3).
Step 2: Simulate the next state (state_2) at time step 2. [instructions]
Step 3: Calculate collaborative actions for state_2.
Step 4: Simulate the next state (state_3) at time step 3. [instructions]
Step 5: Feasibility check for both state_2 and state_3
[task_constraints]
- If any constraint is violated, go back to Step 0 and pick a more conservative action. Repeat Steps 0-4 up to 10 times.
- If no feasible proposal is found after 10 iterations, output the best proposal found.
Output Format (STRICT):
<analysis>[analysis_and_calculation_steps] </analysis>
<proposal>[semantic_proposal_for_negotiation] </proposal>
<output>E[timestep]_R[round]_[agent_name]_proposal = {[my_name]: <proposed action for itself>, [neighbor’s_name]: <proposed acceleration for neighbor>}</output>
IMPORTANT: [self-checking_rules]
Figure 8: The core prompt of CoProposer.
Q.2 Negotiator
Fig. 9 provides the core prompts for proposal evaluation with spatial strategies and feature updates in Negotiator when consensus is not achieved. Fig. 10 provides the core prompts for updating proposal with updated spatial strategies and features in Negotiator.
Refer to caption Negotiator: Proposal Evaluation Query You are [agent_name] for [task_name]. Your observable neighbor is [observable_agents]. In the last negotiation round, consensus was not reached. You need to analyze the disagreement and summarize your reasoning in the following steps:
Step 0: Analyze whether the strategy proposed by [agent_name] conflicts with the strategy proposed by its observable neighbors [agent_name]
[task_strategy_information]
Step 1: According to [Spatial Strategies] and [Spatial features], analyze whether the strategy proposed by [neighbors_agent] conflicts with the trends of [unobservable_agents].
[task_strategy_information]
Output as following format:
<deal>False</deal>
<reasons> Observable vehicles:[observable_agent_name] choose ..., and proposes .... It’s strategy is match/conflict/not conflict with my strategy. The reason is ... I suggest it’s weight in the range ...
[unobservable_agent_list]: They are strategy trend is match/conflict/not conflict with my strategy. The reason is ... I suggest it’s weight in the range ...</reasons>
<spatial-strategy>[update_spatial_strategy]</spatial-strategy>
IMPORTANT: [self-checking_rules]
Figure 9: The core prompt of proposal evaluation and spatial strategy update.
Refer to caption CoProposer: Update Proposal You are [agent_name]. This is the Round_[current_round] of [time_step]. Your observable states is [agent_name Observable State in time_step]. You need to update your proposal by following steps.
- Case 1: If the flag between <deal></deal> is True, reuse the last proposal and output as following format:
<analysis>(I will keep my last proposal, because ...)</analysis>
<proposal> [semantic_proposal_for_negotiation] </proposal>
<output>[structured_proposal_for_coding]</output>
Skip the following steps.
- Case 2: If the flag between <deal></deal> is False:
Step 1: Define the confidence weights for observable neighborsṕroposal and unobservable vehicles weighted average acceleration. [weight_rules]
Output the confidence weights as following format:
<insights>neighbor_proposed_weight:<neighbor_proposed_weight value> unobservable_weight: <unobservable_weight value> my_weight: <my_weight value></insights>
Step 2: Based on weights defined between <insights></insights>, weighted the proposed acceleration for each vehicles by following equations:
updated_acceleration = neighbor_proposed_weight * neighbor_proposed_acceleration + my_weight * my_acceleration + unobservable_weight * unobservable_trend.
/* Step 3: A rollout-simulated verification similar to Step 2-5 in CoProposer */
Output Format (STRICT):
<analysis>[analysis_and_calculation_steps] </analysis>
<proposal>[semantic_proposal_for_negotiation] </proposal>
<output>E[timestep]_R[round]_[agent_name]_proposal = {[agent_name]: <proposed action for itself>,
[neighbor’s_name]: <proposed acceleration for neighbor>}</output>
IMPORTANT: [self-checking_rules]
Figure 10: The core prompt of update proposal.
Q.3 Introspector
Fig. 11 presents the core prompts for generating the revision signal after completing one-step execution. Fig. 12 presents the core prompts for updating the temporal strategy using the drift intensity and semantic revision signal.
Refer to caption Introspector: Generate Semantic Revision Signal You are [agent_name]. You can observe [observable_agent_name]. The total system reward for this time step is xx% lower than in the last time step indicating that the system state getting worse. Your individual reward has dropped from $x$ to $x$, which drop xx%.
Step1. Based on [last_negotiation_record] and [this_negotiation_record], integrate the [Objective Instruction] to analyze whether the drop in episode reward was caused by yourself. If so, at which stage was responsible.
Infer whether the reward decline was caused by [observable_agents] or [unobservable_agents], and output between tags <reasons> </reasons>
Based on <reasons> </reasons> should strategies be adjusted to improve future episode rewards?
Step 1: You need to find out the rules that need to be adjusted based on the reasons you summarized.
Step 2: According to the related rules you summarized, be specific about ‘‘what to do’’ and ‘‘what to avoid’’ and output between tags <self-checking_rules> and </self-checking_rules>.
The output format as follows:
<reasons>[summarized_reasons]</reasons>
<self-checking_rules>(When states condition ... I need to/I need to avoid ...)</self-checking_rules>
IMPORTANT: [self-checking_rules]
Figure 11: The core prompt of generating semantic revision signal.
Refer to caption CoProposer: Update Strategy Query You are [agent_name] to update its strategies. The total system reward for this time step is xx% lower than last time step, indicating that the system state getting worse. The overlap of transition (including state and action) among these two time steps is xx%, which means that their overall movement and decision patterns are xx% similar, with xx% differences in how the agent moves between states and selects actions.
Based on the <reasons></reasons> you summarized (from Fig. 11), update your strategies carefully with NO MORE than xx%.
Step 1. Calculate the change in state from the historical states to the current state and compare it to the previous temporal plan. Analyze if the change in state matches the plan of the previous temporal plan.
Step 2. Carefully check whether the change match the last temporal plan.
- If it matches, keep the original plan and update the temporal plan. Execute step 3 to return the updated plan.
- if it doesn’t match, summarize the reasons and output within <reflect> tags as following: <reflect>(My final action does not match my original temporal plan, the reasons may be ...)</reflect>
According to the <reflect></reflect>, update the temporal plan.
Step 3. Output the updated plan with following format:
<temporal-strategy>[update_temporal_analysis]</temporal-strategy>
IMPORTANT: [self-checking_rules]
Figure 12: The core prompt of update strategy
[self-checking_rules] are a set of rules for verifying the output format and common-sense consistency of LLM outputs. These rules ensure that the model correctly applies physical reasoning. For example, in CPP task, “If ACCELERATION_GAP > 0, MY_VELOCITY will be INCREASED, MY_HEADWAY will be DECREASED.” These rules help correct occasional LLM mistakes, such as misjudging the travel direction of vehicles and thus misunderstanding how acceleration or deceleration affects inter-vehicle headway.
Appendix R AI Assistants in Writing
We utilized AI assistants (e.g., Gemini and Claude) exclusively for grammatical error correction, sentence polishing to enhance readability, and generating robot icons used in Figs. 1 and 2. We explicitly state that the core scientific concepts, experimental design, logical structure, reference, and substantive content of the paper were entirely human-authored without generative contributions from LLMs.